Country XYZ has the following Aggregate Demand Equation:

Y = C + I + G

C = 1,500 + 0.75Y

I = 2,500

G = 6,000

a) Calculate equalibrium Y

b) What is the value of Consumption?

c) What is the value of Marginal Propensity to Consume?
Answer #1

Part a)

Y = C + I + G

Y = 1,500 + 0.75Y + 2,500 + 6,000

0.25Y =10000

Y=10000/0.25 = 40000

Part b)

C = 1,500 + 0.75Y

C = 1,500 + 0.75 x 40000

C = 1,500 + 30000

C = 31,500

Part C)

Marginal Propensity to Consume (MPC) is the change in consumption due to a change in income. To find the same we need to use differentiation formula:

So for every unit of change in Y, C changes by 0.75Y

We can verify the same by substituting a few values:

Y 0 10 30 80
C 1500+0.75 x 0 = 1500 1500+0.75 x10=1507.5 1500+0.75x30=1522.5 1500+0.75x80=1560
Change in Y 10-0=10 30-10=20 80-30=50

Change in C

0.75 x 10 = 7.5

or

1507.5-1500=7.5

0.75x20= 15

or

1522.5-1507.5=15

0.75 x 50=37.5

or

1560-1522.5=37.5